A store contains 1 pair of boots with each of the following colors are black, chocolate and yellow. Each pair is put together in a particular place. You enter into the dark store and pick randomly the boot without looking at it. Then, you replace it with another boots. What is the probability that you will choose the black pair of boots both times?


1
Expert's answer
2022-06-03T11:51:58-0400
"P(BB)=\dfrac{1}{3}(\dfrac{1}{3})=\dfrac{1}{9}"